Circular No. 8793

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 P1 (McNAUGHT) K. Battams, Interferometrics, Inc., and Naval Research Laboratory,

reports that images of comet C/2006 P1 obtained on Jan. 11 with the SECCHI/HI-1B instrument on the NASA STEREO-B (Behind) spacecraft are available at the following website URL: http://ares.nrl.navy.mil/

sungrazer/index.php?p=latest news. The images show a curved, stri-

ated dust tail some 7^{o} long, the bright coma being heavily saturated due to its brightness.

J. N. Marcus writes that forward-scattering brightness enhancement should peak in C/2006 P1, as viewed from the earth, on Jan. 14 at a level ≈ 2.3 magnitudes brighter than what would be predicted by a standard power law (model details in Marcus 2007, ICQ, submitted).

Selected visual total-magnitude, coma-diameter, and tail-length/p.a. estimates made with the comet at low altitude (all magnitudes were corrected by the observers for atmospheric extinction) and in nautical or civil twilight unless otherwise noted: 2006 Dec. 29.28 UT, 3.9, 1′.5, – (B. H. Granslo, Fjellhamar, Norway, 0.10-m refractor); 2007 Jan. 2.30, 2.7, 1′.5, 0^{o}.1 in p.a. 0^{o} (J. J. Gonzalez, Leon, Spain, 25×100 binoculars); 3.28, 1.5, 1′.7, 0^{o}.2 (H. Dahle, Fjellhamar, Norway, 9×63 binoculars); 4.64, 1.0:, –, – (Dahle, Blindern/Oslo, Norway, naked eye); 5.72, −0.5:, –, – (D. Moore, Dublin, Ireland, naked eye); 6.28, 0.2:, 1′, 1^{o} in p.a. 0^{o} (Granslo, Tryvann/Oslo, Norway, 7×50 binoc.); 7.30, 0.0:, 1′, 0^{o}.4 (N. Biver, Meudon, France, 7×50 binoc.); 7.48, −0.5:, about 1′, < 1^{o} (D. W. E. Green, Rowley, MA, U.S.A., 7×35 binoc.); 8.25, −1.2:, 3′, 0^{o}.5 (K. Hornoch, Vranov, Czech Republic, naked eye); 8.65, −0.8:, –, 1^{o}.5 (O. Skilbrei, Honefoss, Norway, naked eye); 8.93, −1.5:, about 1′, – (J. E. Bortle, Stormville, NY, U.S.A., naked eye); 9.68, −2.0, –, – (W. Hasubick, Buchloe, Germany, naked eye); 9.71, −2.2:, 1′, 1^{o}.5 (Biver, naked eye); 9.98, −2.9:, –, short (J. N. Marcus, St. Louis, MO, U.S.A., naked eye); 10.30, −2.2:, –, 1^{o}.0 in p.a. 15^{o} (T. Karhula, Vaesteraas, Sweden, naked eye); 10.34, −2.2:, –, 0^{o}.3 in p.a. 15^{o} (S. Yoshida, Toride, Ibaraki, Japan, 10×70 refractor); 10.66, −2.7:, 5′, 2^{o} in p.a. 25^{o} (K. Hornoch, Krasensko, Czech Rep., 10×50 binoculars); 10.68, −3.2, 2′, 2^{o} (T. Scarmato, Calabria, Italy, 7×50 binoc.); 10.83, −2.5:, –, short (R. A. Keen, Mt. Thorodin, CO, U.S.A., 7.6-cm reflector; broad daylight; comparison with Venus); 10.92, −2.5:, about 1′, 1^{o} (Bortle); 10.95, −2.9:, 1′, 1^{o}.6 in p.a. 15^{o} (P. Creed, New London, OH, U.S.A., 10×50 binoc.); 11.34, −2.8, about 3′, 1^{o} (Y. Nagai, Gunma, Japan, 7×35 binoc.); 11.68, −3.5, 3′, about 2^{o} (Scarmato, naked eye); 11.75, −3.5:, –, 1^{o} (A. Pereira, Cabo da Roca, Portugal, naked eye).

2007 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green